Solve the system of equations -2x-2y=−4 and 4x+3y=11 by combining the equations
Found 2 solutions by richwmiller, rothauserc:
Answer by richwmiller(17219)

combining the equations is also known as adding and subtracting equations as well as the elimination method.
1) -2x-2y=−4
2) 4x+3y=11
We notice that if we multiply the first equation by 2 we can then add the new equation 3) and 2) to eliminate x
3) -4x-4y=−8
2) 4x+3y=11
-y=3
y=-3
substitute -3 in 2)
4x+3y=11
4x+3(-3)=11
4x-9=11
4x=20
x=5
check in 1)
-2x-2y=−4
-2(5)-2(-3)=-4
-10+6=-4
-4=-4
ok

Answer by rothauserc(4718)

1) -2x-2y=−4
:
2) 4x+3y=11
:
multiply equation 1 by 2
:
3) -4x -4y = -8
:
add equation 3 to equation 2
:
-y = 3
:
y = -3
:
substitute for y in equation 1
:
-2x -2(-3) = -4
:
-2x +6 = -4
:
-2x = -10
:
x = 5
:
********************************************
x = 5 and y = -3
:
substitute for x and y in equation 2
:
4(5) +3(-3) = 11
:
20 -9 = 11
:
11 = 11
:
our answer checks
*********************************************
: